OpenStudy (anonymous):

The graph y=(x+2)e^(x^2/2) has one horizontal tangent line. What is the equation of this tangent line?

OpenStudy (anonymous):

having one horizontal tangent line means has a point of which the derivative is zero. y'=e^(x^2/2)+x(x+2)e^(x^2/2)=(x^2+2x+1)e^(x^2/2)=0 since e^(x^2/2)>0 hence x^2+2x+1=0 x=-1 then y=e^(1/2) the equation is Y=e^(1/2) since it is the horizontal line.